27) A certain airplane has a speed of 332.6 km/h and is diving at an angle of 0 = 26.0° below the horizontal when the pilot relea
Answer #1

sol speed of Given at the time of release , the roder decoy is v= 332.6 km/h - 332-645 m/s = 92.388 m/s The angle o= 26.00 Th(b) Since , Let the height of the release point be u for vertical motion, h= (vsind)t + 2 gth distance naut 28th h = (92.388